### v1

@@ Abstract.
Time series forecasting is crucial in various fields such as economics, traffic, and AIOps. However, in real-world applications, focusing solely on the endogenous variables (i.e., target variables), is often insufficient to ensure accurate predictions. Considering exogenous variables (i.e., covariates) provides additional predictive information, thereby improving forecasting accuracy. However, existing methods for time series forecasting with exogenous variables (TSF-X) have the following shortcomings: 1) they do not leverage future exogenous variables, 2) they fail to account for the causal relationships between endogenous and exogenous variables. As a result, their performance is suboptimal. In this study, to better leverage exogenous variables, especially future exogenous variable, we propose a general framework DAG, which utilizes Dual cAusal network along both the temporal and channel dimensions for time series forecasting with exoGenous variables. Specifically, we first introduce the Temporal Causal Module, which includes a causal discovery module to capture how historical exogenous variables affect future exogenous variables. Following this, we construct a causal injection module that incorporates the discovered causal relationships into the process of forecasting future endogenous variables based on historical endogenous variables. Next, we propose the Channel Causal Module, which follows a similar design principle. It features a causal discovery module models how historical exogenous variables influence historical endogenous variables, and a causal injection module incorporates the discovered relationships to enhance the prediction of future endogenous variables based on future exogenous variables. Extensive experiments on multiple datasets demonstrate the state-of-the-art performance of DAG. Additionally, all datasets and code are avaliable at https://github.com/decisionintelligence/DAG.
deep learning; time series forecasting; exogenous variables
journalyear: 2025copyright: acmlicensedconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Time series forecasting plays a vital role in many real-world applications, including economic (Sezer et al., 2020; Huang et al., 2022; Qiu et al., 2025b), traffic (Xu et al., 2023; Mo et al., 2022; Lin et al., 2024a; Sun et al., 2023; Wan et al., 2022), health (Wei et al., 2022; Miao et al., 2021; Lee et al., 2017; Tran et al., 2019), energy (Alvarez et al., 2011; Guo et al., 2015), and AIOps (Wu et al., 2025b; Wang et al., 2023; Qi et al., 2023; Campos et al., 2024; Kieu et al., 2022; Qiu et al., 2025c). Given historical observations, it is valuable if we can know the future values ahead of time. With the rapid development of deep learning, numerous forecasting methods have been proposed (Wu et al., 2025a; Liu et al., 2025a, b, c; Qiu et al., 2025a; Dai et al., 2024; Wu et al., 2024), most of which focus on univariate or multivariate time series and rely on learning the temporal dependencies within a single endogenous (i.e., target) variable or across multiple endogenous variables—see Figure 1a.
However, in addition to the endogenous variables themselves, many practical scenarios involve another type of information that significantly influences the accuracy of predictions—exogenous variables (i.e., covariates). Particularly in scenarios where future exogenous variables are available, effectively leveraging such auxiliary information has the potential to greatly enhance forecasting performance. For example, in traffic prediction, having access to future weather conditions can improve the estimation of future traffic volume; in retail sales forecasting, incorporating holiday schedules and promotional campaigns can lead to better inventory planning. Although these covariates are not part of the prediction target itself, they often contain strong predictive signals.
@@ 1. Introduction
Figure 1. Time series forecasting algorithms can be classified as follows: (a) univariate/multivariate algorithms without exogenous variables, e.g., PatchTST (Nie et al., 2023) and DUET (Qiu et al., 2025d); (b) algorithms considering only historical exogenous variables, e.g., TimeXer (Wang et al., 2024) and CrossLinear (Zhou et al., 2025b); (c) algorithms that account for both historical and future exogenous variables, e.g., TiDE (Das et al., 2023) and TFT (Lim et al., 2021); and (d) algorithms that consider both exogenous variables and causal relationships.
@@
From a temporal perspective, covariates can be categorized into two types: historical exogenous variables, which are exogenous information observed during the historical period, and future exogenous variables, which are exogenous information available at the forecasting time for future time steps. Despite their high predictive value, current deep learning approaches still underutilize future exogenous variables. As shown in Figure 1, existing exogenous-aware forecasting methods can be broadly divided into the following two categories: 1) Methods using only historical information (Figure 1b): These approaches rely solely on historical endogenous and exogenous variables to forecast future endogenous variables. Representative models include TimeXer (Wang et al., 2024) and CrossLinear (Zhou et al., 2025b). By entirely ignoring future covariates, these methods may underperform in scenarios where such information is available in advance. 2) Methods that use both historical and future exogenous variables (Figure 1c): For instance, TiDE (Das et al., 2023) and TFT (Lim et al., 2021) use historical information and future exogenous variables to forecast future endogenous variables. However, without causal constraints, such approaches are susceptible to spurious correlations.
@@
Figure 2. Diagram of Temporal Causality and Channel Causality. Granger causality is used to measure the correlation between historical data and future data. Pearson correlation is used to measure the correlation between exogenous and endogenous variables.
@@
Further analysis, as illustrated in Figure 2, reveals that forecasting with known future covariates involves causal dependencies across both temporal and channel dimensions. Temporally, the influence of historical exogenous variables on future exogenous variables mirrors the evolution from historical to future endogenous variables. In terms of channels, the interaction patterns between historical exogenous and endogenous variables are often transferable to those between future exogenous and endogenous variables. This dual causality structure remains an underexplored but critical feature that is insufficiently addressed by existing methods.
@@
Based on the analysis of existing work and our observations in real-world data, we propose a general framework, DAG, which utilizes Dual cAusal network along both the temporal and channel dimensions for time series forecasting with exoGenous variables, enabling high-quality predictions of future endogenous variables. Specifically, we first introduce the Temporal Causal Module. Since the influence of historical exogenous variables on future exogenous variables structurally resembles the evolution of historical endogenous variables into future endogenous variables, we design a causal discovery module to capture how historical exogenous variables affect future exogenous variables. We then construct a causal injection module that incorporates the discovered causal relationships into the process of forecasting future endogenous variables based on historical endogenous variables. Next, we propose the Channel Causal Module, which follows a similar design principle, where a causal discovery module models how historical exogenous variables influence historical endogenous variables, and a causal injection module incorporates the discovered relationships to enhance the prediction of future endogenous variables based on future exogenous variables. Finally, we combine the temporal causality loss, channel causality loss, and the forecasting loss of future endogenous variables as the overall loss function, enabling end-to-end optimization of the forecasting task.
Our contributions are summarized as:
@@
  • We propose a general framework called DAG, which improves forecasting accuracy by discovering and injecting causal relationships across both temporal and channel dimensions, fully leveraging exogenous variables.
@@
  • We design a causal discovery module to capture how historical exogenous variables influence future exogenous variables and how they affect historical endogenous variables.
@@
  • We design a causal injection module that integrates the discovered temporal and channel causal relationships into the forecasting process of future endogenous variables.
@@
  • We open-source our own collected TSF-X datasets and conduct extensive experiments on both public and newly released datasets. The results demonstrate that DAG outperforms state-of-the-art methods.
2. Related works
2.1. Univariate and Multivariate Forecasting
@@
Existing time series forecasting models are typically classified into univariate and multivariate forecasting methods based on the number of input and output variables. Univariate time series forecasting models rely solely on the historical values of a single variable to predict future values. Traditional univariate forecasting methods, such as ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), and Theta (Federico Garza, 2022) are classical and widely used techniques. However, these methods still depend on manual feature engineering and model design, which limits their flexibility and automation. With the rapid advancement of deep learning technologies (Lu et al., 2024a, 2023, b; Yang et al., 2024; Zhou et al., 2025a; Li et al., 2025b, a; Wang et al., 2025), methods like N-BEATS (Oreshkin et al., 2020) and DeepAR (Salinas et al., 2020) can automatically learn patterns from historical data, excelling at capturing nonlinear relationships and long-term dependencies. On the other hand, multivariate time series forecasting models use multiple input variables to predict the corresponding output variables. The classic methods include VAR (Godahewa et al., 2021), Random Forests(Breiman, 2001) and LightGBM (Ke et al., 2017). In recent years, with the rise of deep learning, various architectural approaches have gained widespread attention. For instance, Transformer architectures, such as Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Triformer (Cirstea et al., 2022a), and PatchTST (Nie et al., 2023), can more accurately capture the complex relationships between temporal tokens. MLP-based methods, such as SparseTSF (Lin et al., 2024b), CycleNet (Lin et al., 2024c), DUET (Qiu et al., 2025d), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), utilize simpler architectures with fewer parameters but still achieve highly competitive forecasting accuracy. However, all those methods overlook an important practical factor—exogenous (historical or future exogenous). In many real-world scenarios, exogenous data is known or can be approximately known, and utilizing exogenous data can significantly improve the accuracy of predictions.
2.2. Forecasting with Exogenous Variables
@@
Time series forecasting with exogenous variables has been extensively discussed in classical statistical methods. Some statistical methods have been extended to incorporate exogenous variables as part of the input. Methods like ARIMAX (Williams, 2001) and SARIMAX (Vagropoulos et al., 2016) have long utilized exogenous variables to enhance forecasting accuracy. More recently, deep learning approaches have advanced this area: CrossLinear (Zhou et al., 2025b) uses cross-correlation embeddings to capture dependencies between historical endogenous and exogenous variables; NBEATSx (Olivares et al., 2023) extends N-BEATS with dedicated branches to utilize both past and future exogenous inputs; TiDE (Das et al., 2023) employs an MLP-based architecture to integrate static and future covariates by concatenation with endogenous features at each time step. The Temporal Fusion Transformer (TFT) (Lim et al., 2021) integrates historical and current exogenous variables using attention mechanisms. Furthermore, TimeXer(Wang et al., 2024) introduces patch-wise embeddings to flexibly incorporate exogenous covariates without strict temporal alignment. On the other hand, ExoTST (Tayal et al., 2024) utilizes innovative embedding, cross-attention, and cross-temporal fusion within an attention framework to robustly handle time lags and missing data. However, these methods generally rely on relatively simple combinations of historical inputs and future exogenous, without fully modeling the complex interactions among historical endogenous, historical exogenous, and future exogenous variables.
3. Preliminaries
3.1. Definitions
@@ 3.2. Problem Statement
$\hat{Y}^{\text{endo}}=\mathcal{F}_{\theta}(X^{\text{endo}},X^{\text{exo}},Y^{\text{exo}})$
@@
where $\hat{Y}^{\text{endo}}\in\mathbb{R}^{N\times F}$ denotes the predicted future endogenous variables.
4. Methodology
@@
Figure 3. The architecture of DAG. (a) Overview of the DAG framework, which comprises Temporal Causal Modules ($\mathcal{F}_{\theta_{1}}$ and $\mathcal{G}_{\theta_{2}}$) and Channel Causal Modules ($\mathcal{F}_{\theta_{3}}$ and $\mathcal{G}_{\theta_{4}}$). (b) Detailed structure of the Temporal Causal Module. (c) Detailed structure of the Channel Causal Module. (c1) The standard Transformer block (Trmblock). (c2) The Causal Trimblock, which injects learned causality into the Trmblock. (d) The loss function. Note that the Gating, Trimblock, and Causal Trimblock used in (b) and (c) share the same architecture.
4.1. Structure Overview
@@
Figure 3 illustrates the overall architecture of DAG, which effectively improves forecasting accuracy by discovering and injecting causal relationships along both the temporal and channel dimensions, thereby fully leveraging the information contained in exogenous variables. Specifically, we first introduce the Temporal Causal Module. Since the influence of historical exogenous variables on future exogenous variables structurally resembles the evolution of historical endogenous variables into future endogenous variables, we design a causal discovery module to capture how historical exogenous variables affect future exogenous variables. We then construct a causal injection module that incorporates the discovered causal relationships into the process of forecasting future endogenous variables based on historical endogenous variables. Next, we propose the Channel Causal Module, which follows a similar design principle, where a causal discovery module models how historical exogenous variables influence historical endogenous variables, and a causal injection module incorporates the discovered relationships to enhance the prediction of future endogenous variables based on future exogenous variables. Finally, we combine the temporal causal loss, channel causal loss, and the forecasting loss of future endogenous variables as the overall loss function, enabling end-to-end optimization of the forecasting task.
4.2. Temporal Causality Module
@@
As analyzed in Section 1, the influence mechanism of historical exogenous variables on future exogenous variables shares a structural similarity with the evolution process from historical endogenous variables to future endogenous variables. Therefore, the Temporal Causality Module is designed to uncover the causal relationships between historical and future exogenous variables (Temporal Causal Discovery Module), and to inject the discovered causality into the modeling process of forecasting future endogenous variables based on historical endogenous variables (Temporal Causal Injection Module), thereby improving prediction accuracy.
4.2.1. Temporal Causal Discovery

### v2

@@ Abstract
Time series forecasting is essential in various domains. Compared to relying solely on endogenous variables (i.e., target variables), considering exogenous variables (i.e., covariates) provides additional predictive information and often leads to more accurate predictions. However, existing methods for time series forecasting with exogenous variables (TSF-X) have the following shortcomings: 1) they do not leverage future exogenous variables, 2) they fail to fully account for the correlation between endogenous and exogenous variables. In this study, to better leverage exogenous variables, especially future exogenous variables, we propose DAG, which utilizes Dual correlAtion network along both the temporal and channel dimensions for time series forecasting with exoGenous variables. Specifically, we propose two core components: the Temporal Correlation Module and the Channel Correlation Module. Both modules consist of a correlation discovery submodule and a correlation injection submodule. The former is designed to capture the correlation effects of historical exogenous variables on future exogenous variables and on historical endogenous variables, respectively. The latter injects the discovered correlation relationships into the processes of forecasting future endogenous variables based on historical endogenous variables and future exogenous variables.
Machine Learning
1 Introduction
@@
Figure 1: Time series forecasting algorithms can be classified as follows: (a) univariate/multivariate algorithms without exogenous variables, e.g., PatchTST (Nie et al., 2023) and DUET (Qiu et al., 2025); (b) algorithms considering only historical exogenous variables, e.g., TimeXer (Wang et al., 2024) and CrossLinear (Zhou et al., 2025); (c) algorithms that account for both historical and future exogenous variables, e.g., TiDE (Das et al., 2023) and TFT (Lim et al., 2021); and (d) algorithms that consider both exogenous variables and correlation relationships.
Time series forecasting plays a vital role in many real-world applications, including economic (Sezer et al., 2020; Huang et al., 2022), traffic (Xu et al., 2023; Mo et al., 2022), health (Wei et al., 2022; Miao et al., 2021), energy (Alvarez et al., 2011; Guo et al., 2015), and AIOps (Wang et al., 2023; Qi et al., 2023). Given historical observations, it is valuable if we can know the future values ahead of time. With the rapid development of deep learning, numerous forecasting methods have been proposed, most of which focus on univariate or multivariate time series and rely on learning the temporal dependencies within a single endogenous (i.e., target) variable or across multiple endogenous variables—see Figure 1a.
However, beyond endogenous variables themselves, many practical scenarios involve another type of information that significantly influences the accuracy of predictions—exogenous variables (i.e., covariates). Particularly in scenarios where future exogenous variables are available, effectively leveraging such auxiliary information can substantially enhance forecasting performance. For example, in traffic prediction, having access to future weather conditions can improve the estimation of future traffic volume; in retail sales forecasting, incorporating holiday schedules and promotional campaigns can lead to better inventory planning. Although these covariates are not part of the prediction target itself, they often contain strong predictive signals.
@@
From a temporal perspective, covariates can be categorized into two types: historical exogenous variables, which are exogenous information observed during the historical period, and future exogenous variables, which are exogenous information known for the forecast horizon. Despite their high predictive value, current deep learning approaches still underutilize future exogenous variables. As shown in Figure 1, existing exogenous-aware forecasting methods can be broadly divided into the following two categories: 1) Methods using only historical information (Figure 1b): These approaches rely solely on historical endogenous and exogenous variables to forecast future endogenous variables. Representative models include TimeXer (Wang et al., 2024) and CrossLinear (Zhou et al., 2025). By entirely ignoring future covariates, these methods may underperform in scenarios where such information is available in advance. 2) Methods that use both historical and future exogenous variables (Figure 1c): For instance, TiDE (Das et al., 2023) and TFT (Lim et al., 2021) use historical information and future exogenous variables to forecast future endogenous variables. However, without fully correlation constraints, such approaches are susceptible to spurious correlations.
@@
Figure 2: Diagram of Temporal Correlation and Channel Correlation. Granger causality is used to measure the correlation between historical data and future data. Pearson correlation is used to measure the correlation between exogenous and endogenous variables.
@@
Further analysis—see Figure 2, reveals that forecasting with known future covariates involves causal dependencies across both temporal and channel dimensions. Temporally, the influence of historical exogenous variables on future exogenous variables mirrors the evolution from historical to future endogenous variables. In terms of channels, the interaction patterns between historical exogenous and endogenous variables are often transferable to those between future exogenous and endogenous variables. This dual causality structure remains an underexplored but critical feature that is insufficiently addressed by existing methods.
@@
Based on the analysis of existing work and our observations in real-world data, we propose a general framework, DAG, which utilizes Dual cAusal network along both the temporal and channel dimensions for time series forecasting with exoGenous variables, enabling high-quality predictions of future endogenous variables. Specifically, we first introduce the Temporal Correlation Module. Since the influence of historical exogenous variables on future exogenous variables structurally resembles the evolution of historical endogenous variables into future endogenous variables, we design a correlation discovery module to capture how historical exogenous variables affect future exogenous variables. We then construct a correlation injection module that incorporates the discovered correlation relationships into the process of forecasting future endogenous variables based on historical endogenous variables. Next, we propose the Channel Correlation Module, which follows a similar design principle, where a correlation discovery module models how historical exogenous variables influence historical endogenous variables, and a correlation injection module incorporates the discovered relationships to enhance the prediction of future endogenous variables based on future exogenous variables. Finally, we combine the temporal correlation loss, channel correlation loss, and the forecasting loss of future endogenous variables as the overall loss function, enabling end-to-end optimization of the forecasting task. Our contributions are summarized as:
@@
  • We propose a general framework called DAG, which improves forecasting accuracy by discovering and injecting correlation relationships across both temporal and channel dimensions, fully leveraging exogenous variables.
@@
  • We design a correlation module to capture historical exogenous impacts on future exogenous and historical endogenous variables.
@@
  • We design a correlation injection module that integrates the discovered temporal and channel correlation into the forecasting process of future endogenous variables.
@@
  • We open-source our own collected TSF-X datasets and conduct extensive experiments on both public and newly released datasets. The results demonstrate that DAG outperforms state-of-the-art methods. Additionally, all datasets and code are available at https://github.com/decisionintelligence/DAG.
2 Related works
2.1 Univariate and Multivariate Forecasting
@@
Existing time series forecasting models are typically classified into univariate and multivariate forecasting methods based on the number of input and output variables. Univariate time series forecasting models rely solely on the historical values of a single variable to predict future values. Traditional univariate forecasting methods, such as ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), and Theta (Garza et al., 2022) are classical and widely used techniques. However, these methods still depend on manual feature engineering and model design, which limits their flexibility and automation. With the rapid advancement of deep learning technologies, methods like N-BEATS and DeepAR can automatically learn patterns from historical data, excelling at capturing nonlinear relationships and long-term dependencies. On the other hand, multivariate time series forecasting models use multiple input variables to predict the corresponding output variables. The classic methods include VAR (Godahewa et al., 2021), Random Forests(Breiman, 2001) and LightGBM (Ke et al., 2017). In recent years, with the rise of deep learning, various architectural approaches have gained widespread attention. For instance, Transformer architectures, such as Informer (Zhou et al., 2021), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Triformer (Cirstea et al., 2022a), and PatchTST (Nie et al., 2023), can more accurately capture the complex relationships between temporal tokens. MLP-based methods, such as SparseTSF (Lin et al., 2024a), CycleNet (Lin et al., 2024b), DUET (Qiu et al., 2025), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), utilize simpler architectures with fewer parameters but still achieve highly competitive forecasting accuracy. However, all those methods overlook an important practical factor—exogenous (historical or future exogenous). In many real-world scenarios, exogenous data is known or can be approximately known, and utilizing exogenous data can significantly improve the accuracy of predictions.
2.2 Forecasting with Exogenous Variables
@@
Time series forecasting with exogenous variables has been extensively discussed in classical statistical methods. Some statistical methods have been extended to incorporate exogenous variables as part of the input. Methods like ARIMAX (Williams, 2001) and SARIMAX (Vagropoulos et al., 2016) have long utilized exogenous variables to enhance forecasting accuracy. More recently, deep learning approaches have advanced this area: CrossLinear (Zhou et al., 2025) uses cross-correlation embeddings to capture dependencies between historical endogenous and exogenous variables; NBEATSx (Olivares et al., 2023) extends N-BEATS with dedicated branches to utilize both past and future exogenous inputs; TiDE (Das et al., 2023) employs an MLP-based architecture to integrate static and future covariates by concatenation with endogenous features at each time step. The Temporal Fusion Transformer (TFT) (Lim et al., 2021) integrates historical and current exogenous variables using attention mechanisms. Furthermore, TimeXer(Wang et al., 2024) introduces patch-wise embeddings to flexibly incorporate exogenous covariates without strict temporal alignment. ExoTST (Tayal et al., 2024) utilizes innovative embedding, cross-attention, and cross-temporal fusion within an attention framework to robustly handle time lags and missing data. On the other hand, GCGNet (Li et al., 2026) models correlations using a graph structure, but does not explicitly distinguish past and future variables or endogenous and exogenous variables. However, these methods generally rely on relatively simple combinations of historical inputs and future exogenous, without fully modeling the complex interactions among historical endogenous, historical exogenous, and future exogenous variables.
3 Preliminaries
3.1 Definitions
@@ 3.2 Problem Statement
$\hat{Y}^{\text{endo}}=\mathcal{F}_{\theta}(X^{\text{endo}},X^{\text{exo}},Y^{\text{exo}}),$ (1)
@@
where $\hat{Y}^{\text{endo}}\in\mathbb{R}^{N\times F}$ denotes the predicted future endogenous variables. $X^{\text{endo}},X^{\text{exo}},Y^{\text{exo}}$ denote historical endogenous variables, historical exogenous variables, and future exogenous variables, respectively.
4 Methodology
@@
Figure 3: The architecture of DAG. (a) Overview of the DAG framework, which comprises Temporal Correlation Modules ($\mathcal{F}_{\theta_{1}}$ and $\mathcal{G}_{\theta_{2}}$) and Channel Correlation Modules ($\mathcal{F}_{\theta_{3}}$ and $\mathcal{G}_{\theta_{4}}$). (b) Detailed structure of the Temporal Correlation Module. (c) Detailed structure of the Channel Correlation Module. (c1) The standard Transformer block (Trmblock). (c2) The Correlation Trmblock, which injects learned correlation into the Trmblock. (d) The loss function. Note that the Gating, Trmblock, and Correlation Trmblock used in (b) and (c) share the same architecture.
4.1 Structure Overview
@@
Figure 3 illustrates the overall architecture of DAG, which effectively improves forecasting accuracy by discovering and injecting correlation relationships along both the temporal and channel dimensions, thereby fully leveraging the information contained in exogenous variables. Specifically, we first introduce the Temporal Correlation Module. Since the influence of historical exogenous variables on future exogenous variables structurally resembles the evolution of historical endogenous variables into future endogenous variables, we design a correlation discovery module to capture how historical exogenous variables affect future exogenous variables. We then construct a correlation injection module that incorporates the discovered correlation relationships into the process of forecasting future endogenous variables based on historical endogenous variables. Next, we propose the Channel Correlation Module, which follows a similar design principle, where a correlation discovery module models how historical exogenous variables influence historical endogenous variables, and a correlation injection module incorporates the discovered relationships to enhance the prediction of future endogenous variables based on future exogenous variables. Finally, we combine the temporal correlation loss, channel correlation loss, and the forecasting loss of future endogenous variables as the overall loss function, enabling end-to-end optimization of the task.
4.2 Temporal Correlation Module
@@
As analyzed in Section 1, the influence mechanism of historical exogenous variables on future exogenous variables shares a structural similarity with the evolution process from historical endogenous variables to future endogenous variables. Therefore, the Temporal Correlation Module is designed to uncover the correlation between historical and future exogenous variables (Temporal Correlation Discovery Module), and to inject the discovered correlation into the modeling process of forecasting future endogenous variables based on historical endogenous variables (Temporal Correlation Injection Module), thereby improving accuracy.
4.2.1 Temporal Correlation Discovery